OpenStudy (anonymous):

What is the graph of the equation? 2x-y=2

OpenStudy (***[isuru]***):

hi, lets make 2x - y = 2 in to a something familiar y = 2x - 2 so this one should be a graph of a straight line which will cut the y axis at -2 and will have a slope of 2 here's the graph http://fooplot.com/plot/7dnjkmxqr3 hope this will help ya!!!

OpenStudy (***[isuru]***):

did ya get it ?

OpenStudy (anonymous):

Yeah. Whattabout this one? "What are the intercepts of the equation? Graph the solution." 5x+3y=15.

OpenStudy (***[isuru]***):

do the same as before... 5x + 3y =15 -----> 3y = 15 - 5x (dividing by 5)-----> y = 3 - x now if the intercept of y = 2x - 2 is -2 what is the intercept of y = -x + 3 ???

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

would it be (3,0) and y-intercept (0,5)?

OpenStudy (***[isuru]***):

no... when an equation is in the form of y = mx + c they say m is the gradient or slope c is the intercept and y = -x + 3 is also in the form of y = mx + c so the intercept is 3 and the graph is.... http://fooplot.com/plot/8wxssgr3qk